OpenStudy (anonymous):

Solve by factoring x^2+7x=-12

OpenStudy (anonymous):

x^2 +7x = -12 x^2 + 7x + 12 = 0 0 = x^2 + 7x + 12; (What adds up to 7 and is also multiplied to equal +12) (Only +3 and +4) = (x + 3)(x+4) Using null factor law x + 3 = 0 x = -3 ------- x +4 = 0 x = -4